# Pushing the frontiers of subexponential FPT time for Feedback Vertex Set

### v1

@@ 2.3. Challenges to lift the result from paths to trees
We now consider the real setting where given $(G,k)$ where $G$ is $K_{r,r}$-free for $r=k^{\varepsilon}$, and given $M$ a feedback vertex set of size at most $2k$, we want to reduce the graph to obtain $|V(G)\setminus M|=\mathcal{O}(k^{2-\varepsilon^{\prime}})$. The approach still consists in partitioning $G-M$ in an appropriate way (called a $t$-uniform partition).
@@
A first problem when trying to adapt the approach of Subsection 2.2 is the degree-related size property. Indeed, after the first two sections Subsection 3.4 and Subsection 3.5, we are now only able to obtain that for any subtree $T$ of $G-M$, $|T|\leq\operatorname{\textsf{poly}}(r)\mu(T)^{\mathcal{O}(\alpha)}$ where $\mu(T)=\max(d_{M}(T),b_{\overline{M}}(T))$ and $b_{\overline{M}}(T)=|\{v\in T,~{}N(v)\not\subseteq M\cup T\}|$ is the size of the “border of $T$”. Observe that $b_{\overline{M}}(P)$ is at most $2$ for any subpath $P$ of path $P_{i}$, whereas $b_{\overline{M}}(T)$ can only be bounded by $|T|$ for a subtree $T$. Informally, in the path case $|P|$ was only polynomially dependent on $d_{M}(P)$, and now $|T|$ is also polynomially depends on $b_{\overline{M}}(T)$.
@@
A second problem is the large degree rule. Suppose that this rule no longer applies (meaning that for every $u\in V(G-M)$, we have $d_{M}(u)\leq t$), and suppose now that because of another rule a vertex $v\in V(G)\setminus M$ is added to $M$, denoting $M^{\prime}=M\cup\{v\}$. Then this can create a new large degree vertex $v^{\prime}$ with $d_{M^{\prime}}(v^{\prime})>t$ (and so $d_{M}(v^{\prime})=t$). Then $v^{\prime}$ would need to be added and the problem may arise again for another vertex $v^{\prime\prime}$. This “cascading” can easily be prevented if $G-M$ is a forest of paths: it suffices to apply the rule a first time at the start of the algorithm, but with $t^{\prime}=t-2$. We then have for each $v\in V(G-M)$ the bound $d_{M}(v)\leq t-2$, and we do not need to applies the rule again after as adding vertices to $M$ may increase $d_{M}(v)$ by at most $2$ ensuring the wanted bound $d_{M}(v)\leq t$ for $v\in V(G-M)$. However, in the case of a tree, we can have in $G-M$ a vertices of arbitrarily large degree, so we cannot apply the same solution. The problem is treated with the help of a technical lemma (that we prove at the end of the proof, see 3.34) which ensures that throughout the execution of the algorithm we keep $|M|=\operatorname{\textsf{poly}}(r)k$.
@@
Finally, a third problem is the definition of the partition. As in the case of paths we want to partition $G-M$ into a “$t$-uniform partition” $\mathcal{T}$, where in particular we have $\mathcal{T}=\mathcal{T}^{+}\cup\mathcal{T}^{-}$, and for any $T\in\mathcal{T}$, $d_{M}(T)\leq 2t$ and $|T|\leq\operatorname{\textsf{poly}}(r)\mu(T)^{\mathcal{O}(\alpha)}$ (see 3.24 for the complete definition). The greedy approach presented for the case of paths is now more involved, as we have to cut each tree of $G-M$ into subtrees that have small border $b_{\overline{M}}(T)$, as otherwise the previous bound $|T|\leq\operatorname{\textsf{poly}}(r)\mu(T)^{\mathcal{O}(\alpha)}$ becomes useless when $\mu(T)$ is too large.
@@
  • Fail and find a subtree $T$ with $|T|>\operatorname{\textsf{poly}}(r)\mu(T)^{C\alpha}$ for some constant $C$, implying that our degree-related size rule can be applied.
@@
The third case is treated in Subsection 3.8, where we either find another way to apply one more time a reduction rule, or prove that $|V(G)\setminus M|=\mathcal{O}(k^{2-\varepsilon^{\prime}})$.
@@ 3.1. Preliminary branching to remove $K_{r,r}$
If $|M_{0}|\leq k_{0}$ or if $|M_{0}|>2k_{0}$ we can conclude that the instance is positive or negative. We thus are left with the case where $k_{0}<|M_{0}|\leq 2k_{0}$.
Let us now describe a branching algorithm starting with $(G_{0},k_{0},M_{0})$ (with $M_{0}$ is a feedback vertex set of $G_{0}$ with $|M_{0}|\leq 2k_{0}$) and leading to a set of instances $\mathcal{I}$, whose properties are discussed below. The algorithm initializes $\mathcal{I}$ to $\{(G_{0},k_{0},M_{0})\}$ and applies the following branching rule to the elements of $\mathcal{I}$ as long as possible.
@@
    Given an instance $(G,k,M)\in\mathcal{I}$, if $G$ contains a $K_{r,r}$-subgraph with parts $A$ and $B$, we replace this instance by $2r$ instances $(G-X,k-(r-1),M\setminus X)$, for any set $X$ of size $r-1$ that is either contained in $A$, or contained in $B$.
Remark 3.2.
If the parameter $k-(r-1)$ is negative, then this instance is negative and we remove it from $\mathcal{I}$.
@@
As $G-M$ is acyclic, for any copy of $K_{r,r}$ with parts $A$ and $B$, one part of this copy is almost fully contained in $M$. More formally and without loss of generality we can assume that $|A\cap M|\geq r-1$. We can thus consider all the $r$ vertices subsets of $V(G)$ consisting of $r-1$ vertices of $M$ and a vertex of $V(G)$ to form the vertex set $A$. Then we can consider each vertex $v$ of $V(G)\setminus A$ and check if $A\subseteq N(v)$. If there are $r$ such vertices, then we have a copy of $K_{r,r}$. It is then trivial to produce the $2r$ instances of (BR1). All this takes $\mathcal{O}(|M|^{r})|V(G)|^{\mathcal{O}(1)}=2^{\mathcal{O}(r\log|M|)}|V(G)|^{\mathcal{O}(1)}$ steps. ∎
We summary the properties obtained after this series of branchings with the following lemma:
@@
    For any $(G,k,M)\in\mathcal{I}$ the graph $G$ is $K_{r,r}$-free induced subgraph of $G_{0}$.
@@
If $(G,k,M)$ is a YES-instance, let $S$ be a feedback vertex set of $G$ with $|S|\leq k$. Any 2 vertices of $A$ induce a 4-cycle with any 2 vertices of $B$, so in order to intersect all cycles $S$ intersects one of $A$ and $B$ on at least $r-1$ vertices. Suppose without loss of generality that $S$ intersects $A$ on (at least) $r-1$ vertices, let $a\in A$ be the $r^{\text{th}}$ vertex of $A$ (i.e. $a$ is a vertex such that $A\setminus\{a\}\subseteq S$, but possibly $a\in S$), and let $X=A\setminus\{a\}$. Let $(G^{\prime},k^{\prime},M^{\prime})$ be the instance, among the $2r$ ones generated by rule (BR1), such that $G^{\prime}=G-X$, and $M^{\prime}=M\setminus X$. Note that taking $S^{\prime}=S\setminus X$, we have that $G-S=G^{\prime}-S^{\prime}$ and $|S^{\prime}|=|S|-(r-1)\leq k-(r-1)=k^{\prime}$. Hence, $(G^{\prime},k^{\prime},M^{\prime})$ is a YES-instance.
Conversely, if $(G^{\prime},k^{\prime},M^{\prime})$ is a YES-instance, with $G^{\prime}=G-X$, and $M^{\prime}=M\setminus X$, let $S^{\prime}$ be a feedback vertex set of $G^{\prime}$ with $|S^{\prime}|\leq k^{\prime}=k-(r-1)$. Note that taking $S=S^{\prime}\cup X$, we have that $G^{\prime}-S^{\prime}=G-S$ and $|S|=|S^{\prime}|+(r-1)\leq k^{\prime}+(r-1)=k$. Hence, $(G,k,M)$ is a YES-instance.
For the second item, if there was a copy of $K_{r,r}$ then the branching process would not be finished. Furthermore, each time rule (BR1) is applied, the new instances are obtained by deleting vertices from the instance formerly in $\mathcal{I}$. Hence, all the generated instances are induced subgraph of the first one, $G_{0}$.
@@
For the last item, let us consider a recursive algorithm $\tilde{A}$ that given $(G,k,M)$ as input, and using (BR1), output all instances of $\mathcal{I}$ generated from $(G,k,M)$ (implying that $\tilde{A}(G_{0},k_{0},M_{0})=\mathcal{I}$). By 3.3, and by observing that for each call of the algorithm we have $|M|\leq 2k_{0}$, the worst case running time $f(n,k)$ of $\tilde{A}$ on a $n$-vertex graph and parameter $k$ is such that $f(n,k)\leq 2^{\mathcal{O}\left(r\log k_{0}\right)}n^{\mathcal{O}(1)}+(2r)f(n,k-(r-1))$. This implies $f(n,k_{0})\leq 2^{\mathcal{O}\left(r\log k_{0}\right)}n^{\mathcal{O}(1)}(2r)^{\frac{k_{0}}{r-1}}$. Moreover, as the number $g(k)$ of instances generated by a call to $\tilde{A}(G,k,M)$ is such that $g(k)\leq(2r)g(k-(r-1))$, we obtain the claimed bound on $|\mathcal{I}|$. ∎
@@ 3.2. The main recursive algorithm
We now consider each element $(G_{i},k_{i},M_{i})$ of $\mathcal{I}$ as an instance $(G_{i},k_{i},M_{i},\emptyset)$ of the following problem $(r,\mathcal{G})$-Ann-FVS, and our goal now is to solve these instances of $(r,\mathcal{G})$-Ann-FVS using our main recursive Algorithm 1.
@@
Given a nice graph class $\mathcal{G}$ and $r>3$, the $(r,\mathcal{G})$-Annotated Feedback Vertex Set problem ($(r,\mathcal{G})$-Ann-FVS for short) is the decision problem where given $(G,k,M,\mathcal{H})$ where $G$ is a $K_{r,r}$-free graph of $\mathcal{G}$, $k$ an integer, $M\subseteq V(G)$ a feedback vertex set of $G$, and $\mathcal{H}$ a family of connected disjoint subsets of $M$ (meaning that for any $H\in\mathcal{H}$, $G[H]$ is connected) such that $|\mathcal{H}|\leq k$, and where the question is whether there exists a feedback vertex set $S$ of $G$ of size at most $k$ that additionally intersects every set of $\mathcal{H}$.
For the sake of completeness we provide here the complete pseudo-code of Algorithm 1, even if it uses rules and subroutine which will be defined later. At this stage, we recommend the reader to only read the following sketch, as the following sections will cover in detail the properties we obtain after each step. The sketch of Algorithm 1 is as follows. We first try to apply (line 2) rules (KR1), (KR2), (KR3), and (KR4), which are like kernelization rules: given the instance $(G,k,M,\mathcal{H})$ we perform a single recursive call on a slightly simpler instance $(G^{\prime},k^{\prime},M^{\prime},\mathcal{H}^{\prime})$. If none of these first rules apply, the algorithm tries to build a special partition of $G-M$ using the Partitionning-Algorithm  . If Partitionning-Algorithm  fails (line 6 or 9) and fall into what we call Case 1 or Case 2, then we apply a kernelization or branching rule. Otherwise we either apply (KR4) or (KR5) (line 16 or line 19), or reach our final point (line 23) where we can prove that $|V(G)|$ is small, implying that $\operatorname{\textsf{tw}}(G)=o(k)$, and solve the instance using a classical DP algorithm.
@@
1:$(G,k,M,\mathcal{H})$ an instance of $(r,\mathcal{G})$-Ann-FVS.
2:if (one of Rule (KR1), Rule (KR2), Rule (KR3), or Rule (KR4) applies on $(G,k,M,\mathcal{H})$then
@@
5:Apply Partitionning-Algorithm  (with $t=2d_{r}$) of 3.27 that tries to build $\mathcal{T}$: a $t$-uniform partition of $G-M$ with $|\mathcal{T}^{+}|\leq p_{3}(r,t)|M|$ (where $p_{3}$ is defined in 3.26)
@@
14:Let $Z_{1}(\mathcal{T})$ and $Z_{2}(\mathcal{T})$ as defined in 3.30, and $\tilde{M}=M\cup Z_{1}(\mathcal{T})\cup Z_{2}(\mathcal{T})$
15:// By 3.31, Rule (KR1) and Rule (KR2) do not apply on $(G,k,\tilde{M},\mathcal{H})$
16:if (Rule (KR4) applies on $(G,k,\tilde{M},\mathcal{H})$, and finds a subtree $T$ that can be removed) then
@@
19:if (Rule (KR5) applies on $(G,k,\tilde{M},\mathcal{H})$ and a connected component $T$ of $G-\tilde{M}$, and finds a vertex $u\in\tilde{M}$ that can be taken) then
@@
22:// $|V(G)|=p_{4}(r,t)|M|$ by 3.32 implying $\operatorname{\textsf{tw}}(G)=o(|M|)$ by Theorem 1.2

### v2

@@ 2.3. Challenges to lift the result from paths to trees
We now consider the general setting where given an instance $(G,k)$, with $G$ being $K_{r,r}$-free for $r=k^{\varepsilon}$, and given $M$ a feedback vertex set of size at most $2k$, we want to reduce the graph to obtain $|V(G-M)|=\mathcal{O}(k^{2-\varepsilon^{\prime}})$. The approach still consists in partitioning GMG-M in an appropriate way (called a $t$-uniform partition).
@@
A first problem when trying to adapt the approach described in Subsection 2.2 above is the degree-related size property. Indeed, after the first two sections Subsection 3.4 and Subsection 3.5, we are now only able to obtain that for any subtree $T$ of GMG-M, $|T|\leq\operatorname{\textsf{poly}}(r)\operatorname{\mu}(T)^{\mathcal{O}(\alpha)}$ where $\operatorname{\mu}(T)=\max(d_{M}(T),b_{\overline{M}}(T))$ and $b_{\overline{M}}(T)=|\{v\in T,~N(v)\not\subseteq M\cup T\}|$ is the size of the “border of TT”. Observe that $b_{\overline{M}}(P)$ is at most $2$ for any subpath $P$ of path $P_{i}$, whereas $b_{\overline{M}}(T)$ can only be bounded by $|T|$ for a subtree $T$. Informally, in the path case $|P|$ could be bounded by a polynomial function of $d_{M}(P)$ only, while now $|T|$ also depends (polynomially) on $b_{\overline{M}}(T)$.
@@
A second problem is the large degree rule. Suppose that this rule no longer applies (meaning that for every $u\in V(G-M)$, we have $d_{M}(u)\leq t$), and suppose now that because of another rule a vertex $v\in V(G-M)$ is added to $M$, denoting $M^{\prime}=M\cup\{v\}$. Then this can create a new large degree vertex $v^{\prime}$ with $d_{M^{\prime}}(v^{\prime})>t$ (and so $d_{M}(v^{\prime})=t$). Then $v^{\prime}$ would need to be added and the problem may arise again for another vertex $v^{\prime\prime}$. This “cascading” could easily be prevented if GMG-M is a union of paths: it suffices to apply the rule a first time at the start of the algorithm, but with $t^{\prime}=t-2$. We then have for each $v\in V(G-M)$ the bound $d_{M}(v)\leq t-2$, and we do not need to apply the rule again after as adding vertices to $M$ may increase $d_{M}(v)$ by at most $2$ ensuring the wanted bound $d_{M}(v)\leq t$ for $v\in V(G-M)$. However, in the case when GMG-M is a forest, it may contain a vertex of arbitrarily large degree, so we cannot apply the same solution. The problem is treated with the help of a technical lemma (that we prove at the end of the proof, see 3.32) which ensures that throughout the execution of the algorithm we keep $|M|=\operatorname{\textsf{poly}}(r)k$.
@@
Finally, a third problem is the definition of the partition. As in the case of paths we want to partition GMG-M into a “$t$-uniform partition” $\mathcal{T}$, where in particular we have $\mathcal{T}=\mathcal{T}^{+}\cup\mathcal{T}^{-}$, and for any $T\in\mathcal{T}$, $d_{M}(T)\leq 2t$ and $|T|\leq\operatorname{\textsf{poly}}(r)\operatorname{\mu}(T)^{\mathcal{O}(\alpha)}$ (see 3.22 for the complete definition). The greedy approach presented for the case of paths is now more involved, as we have to cut each tree of GMG-M into subtrees that have small border $b_{\overline{M}}(T)$, as otherwise the previous bound $|T|\leq\operatorname{\textsf{poly}}(r)\operatorname{\mu}(T)^{\mathcal{O}(\alpha)}$ becomes useless when $\operatorname{\mu}(T)$ is too large.
@@
  • Fail and find a subtree $T$ with $|T|>\operatorname{\textsf{poly}}(r)\operatorname{\mu}(T)^{C\alpha}$ for some constant $C$, implying that our degree-related size rule can be applied.
@@
The third case is treated in Subsection 3.8, where we either find another way to apply one more time a reduction rule, or prove that $|V(G-M)|=\mathcal{O}(k^{2-\varepsilon^{\prime}})$.
@@ 3.1. Preliminary branching to remove $K_{r,r}$
If $|M_{0}|\leq k_{0}$ or if $|M_{0}|>2k_{0}$ we can immediately conclude that the instance is positive or negative. We thus are left with the case where $k_{0}<|M_{0}|\leq 2k_{0}$.
Let us now describe a branching routine leading to a new set of instances $\mathcal{I}$, whose properties are discussed below. The routine initializes $\mathcal{I}$ to $\{(G_{0},k_{0},M_{0})\}$ and applies the following branching rule to the elements of $\mathcal{I}$ as long as possible.
@@
    Given an instance $(G,k,M)\in\mathcal{I}$, if $G$ contains a $K_{r,r}$-subgraph with parts $A$ and $B$, either k<r1k<r-1 and the instance is negative so we remove it from $\mathcal{I}$, or $k\geq r-1$ and we replace the considered instance by the $2r$ instances of the form $(G-X,k-(r-1),M\setminus X)$, for every set $X$ of size r1r-1 that is either subset of $A$, or subset of $B$.
@@
As GMG-M is acyclic, for any copy of $K_{r,r}$ with parts $A$ and $B$, one part of this copy is almost fully contained in $M$. More formally and without loss of generality we can assume that $|A\cap M|\geq r-1$. Therefore, to iterate over all possible choices of $A$ we can consider all the subsets of $V(G)$ consisting of r1r-1 vertices of $M$ and adding to this a vertex of $V(G)$. For each such choice we can consider each vertex $v$ of $V(G)\setminus A$ and check if $A\subseteq N(v)$. If there are $r$ such vertices, then together with $A$ they for a copy of $K_{r,r}$. It is then trivial to produce the $2r$ instances described in (BR1). All this takes $\mathcal{O}(|M|^{r})\cdot|V(G)|^{\mathcal{O}(1)}=2^{\mathcal{O}(r\log|M|)}|V(G)|^{\mathcal{O}(1)}$ steps. ∎
We summarize the properties obtained after this series of branchings with the following lemma:
@@
    For any $(G,k,M)\in\mathcal{I}$ the graph $G$ is a $K_{r,r}$-free induced subgraph of $G_{0}$.
@@
If $(G,k,M)$ is a Yes-instance, let $S$ be a feedback vertex set of $G$ with $|S|\leq k$. Any 2 vertices of $A$ induce a 4-cycle with any 2 vertices of $B$, so in order to intersect all cycles $S$ intersects one of $A$ and $B$ on at least r1r-1 vertices. Suppose without loss of generality that $S$ intersects $A$ on (at least) r1r-1 vertices, let $a\in A$ be the $r^{\text{th}}$ vertex of $A$ (i.e. $a$ is a vertex such that $A\setminus\{a\}\subseteq S$, but possibly $a\in S$), and let $X=A\setminus\{a\}$. Let $(G^{\prime},k^{\prime},M^{\prime})$ be the instance, among the $2r$ ones generated by rule (BR1), such that $G^{\prime}=G-X$, and $M^{\prime}=M\setminus X$. Note that taking $S^{\prime}=S\setminus X$, we have that $G-S=G^{\prime}-S^{\prime}$ and $|S^{\prime}|=|S|-(r-1)\leq k-(r-1)=k^{\prime}$. Hence, $(G^{\prime},k^{\prime},M^{\prime})$ is a Yes-instance.
Conversely, if an instance $(G^{\prime},k^{\prime},M^{\prime})$ generated by (BR1) is a Yes-instance, with $G^{\prime}=G-X$, and $M^{\prime}=M\setminus X$, let $S^{\prime}$ be a feedback vertex set of $G^{\prime}$ with $|S^{\prime}|\leq k^{\prime}=k-(r-1)$. Note that taking $S=S^{\prime}\cup X$, we have that $G^{\prime}-S^{\prime}=G-S$ and $|S|=|S^{\prime}|+(r-1)\leq k^{\prime}+(r-1)=k$. Hence, $(G,k,M)$ is a Yes-instance.
For the second item, if there was a copy of $K_{r,r}$ then the branching process would not be finished. Furthermore, each time rule (BR1) is applied, the new instances are obtained by deleting vertices from the instance formerly in $\mathcal{I}$. Hence, all the generated instances are induced subgraphs of the first one, $G_{0}$.
@@
For the last item, let us consider a recursive algorithm $\tilde{A}$ that given $(G,k,M)$ as input, and using (BR1), output all instances of $\mathcal{I}$ generated from $(G,k,M)$ (implying that $\tilde{A}(G_{0},k_{0},M_{0})=\mathcal{I}$). By 3.2, and by observing that for each call of the algorithm we have $|M|\leq 2k_{0}$, the worst case running time $f(n,k)$ of $\tilde{A}$ on a $n$-vertex graph and parameter $k$ is such that $f(n,k)\leq 2^{\mathcal{O}\left(r\log k_{0}\right)}n^{\mathcal{O}(1)}+(2r)f(n,k-(r-1))$. This implies $f(n,k)\leq 2^{\mathcal{O}\left(r\log k\right)}n^{\mathcal{O}(1)}(2r)^{\frac{k}{r-1}}$. Moreover, as the number $g(k)$ of instances generated by a call to $\tilde{A}(G,k,M)$ is such that $g(k)\leq(2r)g(k-(r-1))$, we obtain the claimed bound on $|\mathcal{I}|$. ∎
@@ 3.2. The main recursive algorithm
We now consider each element $(G_{i},k_{i},M_{i})$ of $\mathcal{I}$ as an instance $(G_{i},k_{i},M_{i},\emptyset)$ of the following auxiliary problem $(r,\mathcal{G})$-Ann-FVS, and our goal now is to solve these instances of $(r,\mathcal{G})$-Ann-FVS using the recursive Algorithm 1 that we described hereafter.
@@
Given a nice graph class $\mathcal{G}$ and an integer $r>3$, the $(r,\mathcal{G})$-Annotated Feedback Vertex Set problem ($(r,\mathcal{G})$-Ann-FVS for short) is the decision problem that takes as input a 4-tuple $(G,k,M,\mathcal{H})$ where:
  • $G$ is a $K_{r,r}$-free graph of $\mathcal{G}$,
  • $k$ an integer,
  • $M\subseteq V(G)$ a feedback vertex set of $G$, and
  • $\mathcal{H}$ a family of at most $k$ connected [^7] That is, for any $H\in\mathcal{H}$, $G[H]$ is connected. disjoint subsets of $M$,
and where the question is to decide whether there is a feedback vertex set $S$ of $G$ of size at most $k$ that additionally intersects every set of $\mathcal{H}$.
For the sake of completeness we provide here the complete pseudo-code of Algorithm 1, even if it uses rules and routines which will be defined later. At this stage, we recommend the reader to only read the following sketch, as the following sections will cover in detail the properties we obtain after each step. The sketch of Algorithm 1 is as follows. We first try to apply (line 2) rules (KR1), (KR2), (KR3), and (KR4), which are like kernelization rules: given the instance $(G,k,M,\mathcal{H})$ we perform a single recursive call on a slightly simpler instance $(G^{\prime},k^{\prime},M^{\prime},\mathcal{H}^{\prime})$. If none of these first rules apply, the algorithm tries to build a special partition of GMG-M using the routine Partitionning-Algorithm  (defined in 3.25). If Partitionning-Algorithm  fails (line 6 or 9) and fall into what we call Case 1 or Case 2, then we apply a kernelization or branching rule. Otherwise we either apply (KR4) or (KR5) (line 16 or line 19), or reach our final point (line 23) where we can prove that $|V(G)|$ is small, implying that $\operatorname{\textsf{tw}}(G)=\mathcal{O}(k^{1-\varepsilon^{\prime}})$, and solve the instance using a classical dynamic programming algorithm.
@@
[…]
2:if (one of Rule (KR1), Rule (KR2), Rule (KR3), or Rule (KR4) applies on $(G,k,M,\mathcal{H})$) then
@@
5:Apply Partitionning-Algorithm  (with $t=2d(r)$) of 3.25 that tries to build $\mathcal{T}$: a $t$-uniform partition of GMG-M with $|\mathcal{T}^{+}|\leq p_{3}(r,t)|M|$ (where $p_{3}$ is defined in 3.24)
@@
14:Let $Z_{1}(\mathcal{T})$ and $Z_{2}(\mathcal{T})$ as defined in 3.28, and $\tilde{M}=M\cup Z_{1}(\mathcal{T})\cup Z_{2}(\mathcal{T})$
[…]
@@
[…]
@@
22:// $|V(G)|=p_{4}(r,t)|M|$ by 3.30 implying $\operatorname{\textsf{tw}}(G)=\mathcal{O}(k^{1-\varepsilon^{\prime}})$ by Theorem 1.3